# Hybrid Quantum Recurrent Neural Network For Remaining Useful Life Prediction

### v1

@@ Abstract
Predictive maintenance in aerospace heavily relies on accurate estimation of the remaining useful life of jet engines. In this paper, we introduce a Hybrid Quantum Recurrent Neural Network framework, combining Quantum Long Short-Term Memory layers with classical dense layers for Remaining Useful Life forecasting on NASA’s Commercial Modular Aero-Propulsion System Simulation dataset. Each Quantum Long Short-Term Memory gate replaces conventional linear transformations with Quantum Depth-Infused circuits, allowing the network to learn high-frequency components more effectively. Experimental results demonstrate that, despite having fewer trainable parameters, the Hybrid Quantum Recurrent Neural Network achieves up to a 5% improvement over a Recurrent Neural Network based on stacked Long Short-Term Memory layers in terms of mean root mean squared error and mean absolute error. Moreover, a thorough comparison of our method with established techniques, including Random Forest, Convolutional Neural Network, and Multilayer Perceptron, demonstrates that our approach, which achieves a Root Mean Squared Error of 15.46, surpasses these baselines by approximately 13.68%, 16.21%, and 7.87%, respectively. Nevertheless, it remains outperformed by certain advanced joint architectures. Our findings highlight the potential of hybrid quantum-classical approaches for robust time-series forecasting under limited data conditions, offering new avenues for enhancing reliability in predictive maintenance tasks.
@@ I Introduction
Accurate estimation of the remaining useful life (RUL) of critical machinery is a cornerstone of modern reliability and risk-management strategies si2011remaining ; lee2014prognostics ; berghout2022systematic . Nowhere is this more evident than in the aviation industry: the timely prediction of gas turbine jet engine failures not only bolsters safety by preventing catastrophic breakdowns zhang2016multiobjective , but also serves to streamline maintenance and resource allocation kang2021remaining . Indeed, robust forecasting models enable operators to defer costly repairs until necessary while avoiding the risks of overdue overhauls, ultimately leading to reduced downtime and more efficient fleet utilization huang2019bidirectional ; ferreira2022remaining .
@@
Finally, the new tool that allows the use of the advantages of ML algorithms and quantum computations emerged — quantum machine learning (QML), which proved its usability by solving problems typified by constrained datasets or high-dimensional feature spaces rebentrost2014quantum ; ciliberto2018quantum ; schuld2018supervised ; cao2017quantum ; Alharbi2024QCNN . QML utilises high-dimensional Hilbert spaces to encode input features, allowing it to detect more complex patterns with fewer parameters than its classical counterpart havlivcek2019supervised ; schuld2019quantum . This proved to be especially useful when applied to problems with nonstationary or noisy signals, which are frequently encountered in industrial maintenance and prognostic applications  sagingalieva2023photovoltaic ; emmanoulopoulos2022quantummachinelearningfinance ; protasevich2025multi ; lee2025predictive .
@@
A promising approach for leveraging these quantum advantages involves hybrid quantum-classical neural networks (HQNNs), which combine classical and quantum ML in one single framework arthur2022hybridquantumclassicalneuralnetwork ; sagingalieva2023photovoltaic ; sagingalieva2023hybrid ; Bischof2025Hybrid ; SUN2025130226 . Inside this structure, the quantum component plays a specific role in data processing or feature encoding while leaving computationally hard tasks like backpropagation and parameter optimization for the classical routines. By selectively introducing quantum effects, HQNNs retain the scalability and reliability of conventional deep learning pipelines broughton2021tensorflowquantumsoftwareframework . Preliminary studies indicate that HQNNs can match or exceed the performance of their classical counterparts, often exhibiting greater resilience against overfitting abbas2021power ; Berberich2024training . The inherent parameter efficiency of HQNNs further enhances their suitability for scenarios involving complex, high-dimensional inputs or limited data availability — situations frequently observed in predictive maintenance sagingalieva2023hybrid ; anoshin2024hybrid .
@@
In this paper, we propose a novel Hybrid Quantum Recurrent Neural Network (HQRNN) tailored for jet engine RUL forecasting. Our approach replaces the linear transformations in each gate of a classical LSTM with a Quantum Depth-Infused (QDI) circuit sagingalieva2023hybrid ; anoshin2024hybrid ; lusnig2024hybrid , creating a quantum-enhanced LSTM (QLSTM) chen2022quantum ; sagingalieva2023photovoltaic . By encoding sensor measurements in a quantum Hilbert space, these QDI layers effectively capture complex temporal patterns while requiring fewer trainable parameters than a purely classical baseline chen2022quantum . We evaluate our method on the NASA Commercial Modular Aero-Propulsion System Simulation (C-MAPSS) dataset saxena2008turbofan , a widely used benchmark in turbofan engine degradation studies, and demonstrate improved predictive performance in terms of both root mean squared error (RMSE) and mean absolute error (MAE). Moreover, experimental results suggest that our hybrid approach can help mitigate the challenges associated with limited data, underscoring its potential benefits in industrial time-series applications sagingalieva2023hybrid .
@@
The C-MAPSS dataset is a widely used benchmark for predicting RUL of aircraft engines. Recent studies addressing this prediction problem include approaches based on Temporal Convolutional Neural Networks (TCNN) combined with Transformers wang2021remaining and joint CNN-LSTM models enhanced with attention mechanisms deng2024prediction ; both articles also provide comprehensive summaries of alternative methods. In Section VI, we present a detailed comparative analysis of our proposed HQRNN against classical machine learning techniques and simple neural network architectures (Table 2), as well as recent state-of-the-art joint deep learning approaches (Table 3). Our HQRNN approach demonstrates superior performance compared to traditional ML and simple neural network models, though it currently underperforms relative to several advanced joint models. These results highlight promising opportunities for future research focused on integrating quantum-enhanced recurrent structures with complementary deep learning methods.
@@ II Dataset
The NASA C-MAPSS dataset is a widely used multivariate time series for analyzing gas turbine engine degradation. In this work, we employ the “FD001” subset, which includes run-to-failure data for a fleet of 100 engines. Each engine progresses from a nominal state to the point of failure, providing sensor measurements over multiple cycles. The dataset comprises four main feature groups:
@@
The data are split into a training set and a test set. In the training set, each engine’s sensor measurements are recorded until failure occurs, whereas in the test set, the measurements end at an arbitrary cycle prior to failure. So the goal is to predict the number of remaining operational cycles before failure in the test set, i.e., the number of operational cycles after the last cycle that the engine will continue to operate. A vector of true RUL values for the test data is provided.
We design our predictive model as follows. A model gets a fixed size window of consecutive engine cycles as input and yields a single estimated RUL value as output. This means that for each entry in the training dataset there should be a corresponding target RUL value.
Creation of these RUL values depends on the degradation model. A linear degradation model, for instance, linearly decreases the RUL value to zero as engine approaches the point of failure. In this work, however, we will use a piecewise linear degradation model, which follows the idea that for some time an engine is operating normally (RUL is constant, also known as “early RUL”) and at some point develops a fault, which leads to linear decrease of its useful life, see Figure 1. We chose to set early RUL to 125 cycles, based on the minimum run length of 128 cycles observed in the training data.
@@ III Hybrid Quantum Recurrent Neural Network
Figure 2 provides an overview of the HQRNN model and its constituent layers. Fig. 2(a) shows the complete pipeline: an input window of sensor measurements of size $W\times 14$ is passed through three stacked QLSTM layers with latent dimensions of 32, 16, and 8, respectively. The output of the QLSTM stack is then fed into classical Dense layers, which reduce the dimension from $8\times W$ to 16, then 32, and finally to a single value representing the predicted RUL. Fig. 2(b) depicts the internal structure of a single QLSTM layer, highlighting the replacement of each conventional linear gate transformation with a individual QDI layer.
@@
[…]
$\text{CNOT}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\\ \end{pmatrix}$
@@ IV Training and Results
Using an Adam optimizer with a batch size of 128 and learning rate $0.001$, we train both the HQRNN and a purely classical RNN for 20 epochs. The two models have an identical number of parameters to ensure a fair comparison.
Each model is trained with 10 different random seeds, and we take the average of the resulting predictions. Surprisingly, as shown in Table 1, the HQRNN outperforms the classical RNN in both RMSE and MAE across all tested parameter configurations. This observation appears to contradict the training-phase results, yet it is consistent with the premise that quantum models may better generalize in data-scarce scenarios caro2022generalization ; schuld2021effect (remember that there are just a $100$ engines in the training dataset).
@@
Model Mean RMSE Best RMSE Mean MAE Best MAE N of parameters
@@ V.2 Trainability Analysis: Fisher Information
In supervised machine learning, one typically trains a model $h_{\theta}(\hat{x})$ on a labeled dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x$ represents input data and $y$ denotes the corresponding labels (targets). The model parameters $\theta=(\theta^{1},\theta^{2},\dots,\theta^{n})$ define a conditional probability distribution $p(y\mid x,\theta)$, which can also be expressed via the joint distribution $p(x,y\mid\theta)$ as
@@
The set of all such distributions for different $\theta$ forms a manifold =$\mathcal{M}=\{p(y\mid x,\theta),\,\theta\in\Theta\}$.
@@
Figure 4: (a) The normalized histogram of the Fisher eigenvalue spectrum. There are several eigenvalue groups with slightly higher contribution; thus, the circuit doesn’t fully rely on just few significant parameters, especially as all the others follow them closely in terms of eigenvalue percentage. Furthermore, according to abbas2021power , small number of close to zero eigenvalues indicates the resilience to the barren plateau problem. Additionally, none of those groups come close to the $95\%$ majority threshold, thus showing equal contributive distribution. (b) This is an averaged normalized Fisher Information Matrix. The diagonal of this matrix shows that the quantum circuit equally distributes the gradients to all trainable parameters, and there is no evident single-parameter dominance. Anti-diagonal elements aren’t pronounced which means parameters aren’t interconnected resulting in easier optimisation for circuit’s weights.
@@ VI Discussion
In this study, we presented a HQRNN for predicting RUL of jet engines. Our model employs stacked QLSTM layers, wherein the conventional linear transformations of each LSTM gate are replaced by QDI circuits. As demonstrated in Section V, these QDI layers exhibit strong expressivity, capturing a broad spectrum of frequencies while maintaining a parameter-efficient design.
We tested the HQRNN on the NASA C-MAPSS “FD001” dataset, along with a classical RNN of comparable parameter size. Although the classical model achieved slightly lower losses on the training and validation sets, the HQRNN demonstrated superior generalization on the test set. Specifically, it showed approximately a 5% improvement in both mean RMSE and MAE while also requiring fewer learnable parameters. This suggests that the quantum-enhanced network may be less prone to overfitting and more adept at capturing diverse frequency components that are critical for RUL prediction, especially when data are limited or particularly complex. Analysis of the quantum circuit via ZX calculus, Fisher Information, and Fourier series further confirmed that the chosen QDI circuit is both trainable and expressively powerful, offering a viable path toward improved modeling of real-world degradation patterns.
In comparison with classical ML models (random forest, LASSO regression, etc) and simple ANN models (MLP, CNN, LSTM) on this task, proposed HQRNN model shows the best performance in terms of RMSE error (Table 2).
However, if one wants to achieve state-of-the-art results in such tasks, it is better to use ensembles or combinations of neural network models with addition of complex feature preprocessing (Table 3). For instance, “Auto RUL + LSTM model” asif2022deep uses an improved sophisticated degradation model to assing RUL target labels, while this work sticks to simple piecewise linear degradation model. So it is expected that joint models represented in Table 3 are performing better than our raw HQRNN model. This means that the proposed model can be used not only on its own, but can also be integrated in complex ML pipelines, to further improve their performance.

### v2

@@ Abstract
Predictive maintenance in aerospace heavily relies on accurate estimation of the remaining useful life of jet engines. In this paper, we introduce a Hybrid Quantum Recurrent Neural Network framework, combining Quantum Long Short-Term Memory layers with classical dense layers for Remaining Useful Life forecasting on NASA’s Commercial Modular Aero-Propulsion System Simulation dataset. Each Quantum Long Short-Term Memory gate replaces conventional linear transformations with Quantum Depth-Infused circuits, allowing the network to learn high-frequency components more effectively. Experimental results demonstrate that, despite having fewer trainable parameters, the Hybrid Quantum Recurrent Neural Network achieves up to a $5\%$ improvement over a Recurrent Neural Network based on stacked Long Short-Term Memory layers in terms of mean root-mean-square error and mean absolute error. Moreover, a thorough comparison of our method with established techniques, including Random Forest, Convolutional Neural Network, and Multilayer Perceptron, demonstrates that our approach, which achieves a Root Mean Squared Error of $15.46$, surpasses these baselines by approximately $13.68\%$, $16.21\%$, and $7.87\%$, respectively. Nevertheless, certain advanced joint architectures still outperform it. Our findings highlight the potential of hybrid quantum-classical approaches for robust time-series forecasting under limited-data conditions, offering new avenues for enhancing reliability in predictive maintenance tasks.
@@ I Introduction
Accurate estimation of the remaining useful life (RUL) of critical machinery is a cornerstone of modern reliability and risk management strategies si2011remaining ; lee2014prognostics ; berghout2022systematic . Nowhere is this more evident than in the aviation industry — the timely prediction of gas turbine jet engine failures not only bolsters safety by preventing catastrophic breakdowns zhang2016multiobjective , but also serves to streamline maintenance and resource allocation kang2021remaining . Indeed, robust forecasting models enable operators to defer costly repairs until necessary while avoiding the risks of overdue overhauls, ultimately leading to reduced downtime and more efficient fleet utilization huang2019bidirectional ; ferreira2022remaining .
@@
Finally, the new tool that allows the use of the advantages of ML algorithms and quantum computations emerged — quantum machine learning (QML), which proved its usability by solving problems typified by constrained datasets or high-dimensional feature spaces rebentrost2014quantum ; ciliberto2018quantum ; schuld2018supervised ; cao2017quantum ; Alharbi2024QCNN ; qml_review_2023 . QML utilities high-dimensional Hilbert spaces to encode input features, allowing it to detect more complex patterns with fewer parameters than its classical counterpart havlivcek2019supervised ; schuld2019quantum . This proved to be especially useful when applied to problems with nonstationary or noisy signals, which are frequently encountered in industrial maintenance and prognostic applications emmanoulopoulos2022quantummachinelearningfinance ; sedykh2024hybrid ; sagingalieva2025photovoltaic .
@@
A promising approach for leveraging these quantum advantages involves hybrid quantum-classical neural networks (HQNNs), which combine classical and quantum ML in one single framework kordzanganeh2023parallel ; kordzanganeh2023exponentially ; arthur2022hybridquantumclassicalneuralnetwork ; haboury2024information ; Bischof2025Hybrid ; SUN2025130226 ; patapovich2025superposed . Inside this structure, the quantum component plays a specific role in data processing or feature encoding while leaving computationally hard tasks like backpropagation and parameter optimization for the classical routines. By selectively introducing quantum effects, HQNNs retain the scalability and reliability of conventional deep learning pipelines broughton2021tensorflowquantumsoftwareframework . Preliminary studies indicate that HQNNs can match or exceed the performance of their classical counterparts sagingalieva2023hyperparameter ; senokosov2024quantum ; haboury2023supervised ; sagingalieva2025hybrid , often exhibiting greater resilience against overfitting abbas2021power ; Berberich2024training . The inherent parameter efficiency of HQNNs further enhances their suitability for scenarios involving complex, high-dimensional inputs or limited data availability — situations frequently observed in predictive maintenance.
@@
In this paper, we propose a novel Hybrid Quantum Recurrent Neural Network (HQRNN) tailored for jet engine RUL forecasting. Our approach replaces the linear transformations in each gate of a classical LSTM with a Quantum Depth-Infused (QDI) circuit sagingalieva2023hybrid ; anoshin2024hybrid ; lusnig2024hybrid , creating a quantum-enhanced LSTM (QLSTM) chen2022quantum ; sagingalieva2025photovoltaic . By encoding sensor measurements in a quantum Hilbert space, these QDI layers effectively capture complex temporal patterns while requiring fewer trainable parameters than a purely classical baseline chen2022quantum . We evaluate our method on the NASA Commercial Modular Aero-Propulsion System Simulation (C-MAPSS) dataset saxena2008turbofan , a widely used benchmark in turbofan engine degradation studies, and demonstrate improved predictive performance in terms of both root mean squared error (RMSE) and mean absolute error (MAE). Moreover, experimental results suggest that our hybrid approach can help mitigate the challenges associated with limited data, underscoring its potential benefits in industrial time-series applications sagingalieva2025photovoltaic ; kurkin2025forecasting ; lee2025predictive .
@@
The C-MAPSS dataset cmapss is a widely used benchmark for predicting RUL of aircraft engines. Recent studies addressing this prediction problem include approaches based on Temporal Convolutional Neural Networks (TCNN) combined with Transformers wang2021remaining and joint CNN-LSTM models enhanced with attention mechanisms deng2024prediction ; both articles also provide comprehensive summaries of alternative methods. In Section VI, we present a detailed comparative analysis of our proposed HQRNN against classical machine learning techniques and simple neural network architectures (Table 2), as well as recent state-of-the-art joint deep learning approaches (Table 3). Our HQRNN approach demonstrates superior performance compared to traditional ML and simple neural network models, though it currently underperforms relative to several advanced joint models. These results highlight promising opportunities for future research focused on integrating quantum-enhanced recurrent structures with complementary deep learning methods.
@@ II Dataset
The NASA C-MAPSS dataset cmapss is a widely used multivariate time series for analyzing gas turbine engine degradation. In this work, we employ the “FD001” subset, which includes run-to-failure data for a fleet of 100 engines. Each engine progresses from a nominal state to the point of failure, providing sensor measurements over multiple cycles. The dataset comprises four main feature groups:
@@
The data are split into a training set and a test set. In the training set, each engine’s sensor measurements are recorded until failure occurs, whereas in the test set, the measurements end at an arbitrary cycle prior to failure. The goal is to predict the number of remaining operational cycles before failure in the test set, i.e., the number of operational cycles after the last cycle that the engine will continue to operate. A vector of true RUL values for the test data is provided.
We design our predictive model as follows. A model gets a fixed-size window of consecutive engine cycles as input and yields a single estimated RUL value as output. This means that for each entry in the training dataset there should be a corresponding target RUL value.
Creation of these RUL values depends on the degradation model. A linear degradation model, for instance, linearly decreases the RUL value to zero as the engine approaches the point of failure. In this work, however, we will use a piecewise linear degradation model, which follows the idea that for some time an engine is operating normally (RUL is constant, also known as “early RUL”) and at some point develops a fault, which leads to linear decrease in its useful life, see Figure 1. We choose to set early RUL to 125 cycles, based on the minimum run length of 128 cycles observed in the training data.
@@ III Hybrid Quantum Recurrent Neural Network
Figure 2 provides an overview of the HQRNN model and its constituent layers. Fig. 2(a) shows the complete pipeline: an input window of sensor measurements of size $W\times 14$ is passed through three stacked QLSTM layers with latent dimensions of 32, 16, and 8, respectively. The output of the QLSTM stack is then fed into classical Dense layers, which reduce the dimension from $8\times W$ to 16, then 32, and finally to a single value representing the predicted RUL. Fig. 2(b) depicts the internal structure of a single QLSTM layer, highlighting the replacement of each conventional linear gate transformation with an individual QDI layer.
@@
$R_{x}(\theta)=\begin{pmatrix}\cos(\frac{\theta}{2})&-i\sin(\frac{\theta}{2})\\[4.0pt] -i\sin(\frac{\theta}{2})&\ \cos(\frac{\theta}{2})\end{pmatrix}$
$\mathrm{CNOT}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\\ \end{pmatrix}$
@@ IV Training and Results
Using an Adam optimizer with a batch size of 128 and a learning rate of $0.001$, we train both the HQRNN and a purely classical RNN for 20 epochs. The two models have an identical number of parameters to ensure a fair comparison.
Each model is trained with 10 different random seeds, and we take the average of the resulting predictions. Surprisingly, as shown in Table 1, the HQRNN outperforms the classical RNN in both RMSE and MAE across all tested parameter configurations. This observation appears to contradict the training-phase results, yet it is consistent with the premise that quantum models may better generalize in data-scarce scenarios caro2022generalization ; schuld2021effect (remember that there are just $100$ engines in the training dataset).
@@
Model Mean RMSE Best RMSE Mean MAE Best MAE Number of parameters
@@ V.2 Trainability Analysis: Fisher Information
[…]
@@
The set of all such distributions for different $\theta$ forms a manifold $\mathcal{M}=\{p(y\mid x,\theta),\,\theta\in\Theta\}$.
@@
Figure 4: (a) The normalized histogram of the Fisher eigenvalue spectrum. There are several eigenvalue groups with slightly higher contribution; thus, the circuit doesn’t fully rely on just few significant parameters, especially as all the others follow them closely in terms of eigenvalue percentage. Furthermore, according to abbas2021power , small number of close to zero eigenvalues indicates the resilience to the barren plateau problem. Additionally, none of those groups come close to the $95\%$ majority threshold, thus showing equal contributive distribution. (b) This is an averaged normalized Fisher Information Matrix. The diagonal of this matrix shows that the quantum circuit equally distributes the gradients to all trainable parameters, and there is no evident single-parameter dominance. Anti-diagonal elements aren’t pronounced which means parameters aren’t interconnected resulting in easier optimization for circuit’s weights.
@@ VI Discussion
In this study, we presented a HQRNN for predicting the RUL of jet engines. Our model employs stacked QLSTM layers, wherein the conventional linear transformations of each LSTM gate are replaced by QDI circuits. As demonstrated in Section V, these QDI layers exhibit strong expressivity, capturing a broad spectrum of frequencies while maintaining a parameter-efficient design.
We tested the HQRNN on the NASA C-MAPSS “FD001” dataset, along with a classical RNN of comparable parameter size. Although the classical model achieved slightly lower losses on the training and validation sets, the HQRNN demonstrated superior generalization on the test set. Specifically, it showed approximately a $5\%$ improvement in both mean RMSE and MAE while also requiring fewer learnable parameters. This suggests that the quantum-enhanced network may be less prone to overfitting and more adept at capturing diverse frequency components that are critical for RUL prediction, especially when data are limited or particularly complex. Analysis of the quantum circuit via ZX calculus, Fisher Information, and Fourier series further confirmed that the chosen QDI circuit is both trainable and expressively powerful, offering a viable path toward improved modeling of real-world degradation patterns.
In comparison with classical ML models (random forest, LASSO regression, etc) and simple ANN models (MLP, CNN, LSTM) on this task, the proposed HQRNN model shows the best performance in terms of RMSE error (Table 2).
However, if one wants to achieve state-of-the-art results in such tasks, it is better to use ensembles or combinations of neural network models with the addition of complex feature preprocessing (Table 3). For instance, “Auto RUL + LSTM model” asif2022deep uses an improved sophisticated degradation model to assign RUL target labels, while this work sticks to simple piecewise linear degradation model. Therefore, it is expected that joint models represented in Table 3 are performing better than our raw HQRNN model. This means that the proposed model can be used not only on its own but also integrated in complex ML pipelines, to further improve their performance.
@@
Conflict of interest
The authors declare no conflicts of interest.
Funding
No funding was received for this manuscript.